OpenStudy (anonymous):

3 1/2+2 1/8+ 1 1/4=?

OpenStudy (anonymous):

i've changed the fractions to improper fractions: 3 1/2+2 1/8+ 1 1/4=? 7/2+17/8+5/4=? you want to get all the fractions under the same denominator, so we'll put them over 8. 7/2=28/8 17/8=17/8 5/4=10/8 28/8+17/8+10/8=? (28+17+10)/8=? 55/8=? ?=6 7/8